# Statistics within UV–Visible Absorption Spectrum of Ethanolic Azobenzene

### v1

@@ Abstract
We proceed with the discussion of the results. The observed strong statistical variation of the UV–visible spectrum highlights the importance of proper ensemble sampling for the analysis of spectra of liquid systems. In this case, the variation is similar, if not greater, to what has been reported for X-ray spectra of liquids [2, 4, 31, 12, 13]. Furthermore, our analysis shows that the spectrum is affected by a wide range of structural properties, as a multitude of features in the ECA curves were non-zero (see Figure 3 and Supplementary Information). In spite of the large number of involved features, we emphasize that this subspace still covers a relatively small fraction of the total structural variance (Figures 1b and 1c). Last, the fact that the ECA curves show a contribution from oxygen atoms is a direct indication that the ethanol solvent has a significant effect on the spectrum.
@@
The presented analysis necessitated numerous approximations. For example, the cutoff value for the explicit solvent is an obvious one. We chose the cutoff to be as large as possible while still enabling the simulation of $\sim$104 structure–spectrum pairs. Additionally, we ignore the possible spectral effects arising from photodynamics and vibrational states. Despite this, the reasonability of our data is supported by the fact that the general shape of the spectrum agrees with the experiment. Regarding the analysis of this work, the challenges and limitations of the ECA method have been discussed in detail in Ref. 13. Most notably, the interpretation of the ECA results is dependent on the properties of the descriptor. Therefore, development of a new, highly interpretable, atomistic descriptor for machine learning can have a significant positive impact on the method in the future.
@@
Instead of representing two specific atomistic structures, the curves in Figure 3 reflect shifts in the structural distributions between the two selected points in the latent coordinate space. While not necessarily always the case, for the UV–visible spectrum of ethanolic trans-azobenzene, the two latent coordinates link the respective structural classes to a shift of the S2 line, and reveal potentially important implications for photoprocesses in liquids. Namely, as a consequence of the observed spectral behaviour, the used excitation wavelength introduces a preference for certain initial structures for the following photodynamics, potentially observable by, e.g., pump–probe experiments. Solvent properties have also been found to have an effect on the photochemistry of azobenzene [15], which could partially be a manifestation of this pre-selection.
@@
In conclusion, the UV–visible absorption spectrum of ethanolic trans-azobenzene shows a broad statistical variation analogous to that found in X-ray spectra. Our subsequent analysis condensed the spectral variance to a few structural degrees of freedom, enabling a partial interpretation of the complex structure–spectrum relationship. We found that photon excitation on the short-wavelength side of the S2 peak favors structures with, among other characteristics, weaker hydrogen bonding with the ethanol solvent, and contraction of the N=N bond. Thus, the chosen incident photon energy may favor certain structures for the following photodynamics.
@@ Acknowledgements
We acknowledge CSC – IT Center for Science, Finland, for providing computing resources. We thank Dr. N. Runeberg at CSC for support with the Puhti supercomputer, and Prof. J. Wachtveitl for providing the experimental UV–visible absorption spectrum of trans-azobenzene in ethanol solution.
@@
.1 Machine learning details
@@
number of grid points {10, 20, 30}
@@
.2 A schematic of all LMBTR centers
@@
.3 ROI value difference with large overall intensity: full list of structural changes
@@
Figure 6 shows the complete list of structural changes implied by a blueshift in the S2 peak, according to the applied analysis procedure. The component vector, originally in the z-score-standardized descriptor space, has been inverse transformed to the original descriptor space after which each feature has been divided by the squared distance to the corresponding center. As mentioned in the main text, several curves show behavior indicative of bending of the N=N–C angles. In addition to the results discussed in the main text, the curves at centers C2 and C13 also show a shortening of the C–C bonds. The virtual centers located off-plane either from the nitrogen atoms or the carbon rings show large values near zero. This is an artifact originating from the division by the squared distance to the center. We note that the curves also show other effects that are difficult to interpret; for example the hydrogen and carbon curves of the virtual centers.
@@
Figure 6: The full LMBTR vector corresponding to the structural shift behind the increase of the ROI value difference. We denote regions corresponding to weakening of hydrogen bonds with an asterisk (*), shortening of N=N bond with triangles ($\triangledown$), shortening of or C–C bonds with diamonds ($\diamond$), increase in the N–C bond lengths with squares ($\Box$), and bending of the N=N–C angles with circles ($\circ$). Numerous other less interpretable factors also play a role in the structure–spectrum relationship.
.4 Structural changes implied by the increase of total ROI value
@@
Similarly to the analysis of the ROI value difference, Figure 7 shows the complete list of structural changes implied by an increase in the total ROI value, i.e. the sum of the average differential cross section within each ROI. Again, several structural factors affect the spectrum, some being more interpretable than others. These include, for example, presence of ethanol solvent near the tips of the azobenzene rings. As is the case in the analysis of the ROI difference, twisting of the azobenzene molecule is also implied. In contrast to the ROI difference, changes in hydrogen bonding between the azobenzene molecule and the ethanol solvent do not have a notable effect on the total ROI value.
@@
Figure 7: The full LMBTR vector corresponding to the structural shift behind the increase in total ROI value. We denote regions corresponding to bending of the N=N–C angles with circles ($\circ$) and increase in the amount of ethanol solvent near the tips of the azobenzene rings with plus signs (+). Many other less interpretable factors also play a role in the structure–spectrum relationship.

### v2

@@ IV Discussion
The observed strong statistical variation of the UV–visible spectrum highlights the importance of proper ensemble sampling for the analysis of spectra of liquid systems. In this case, the variation is similar, if not greater, to what has been reported for X-ray spectra of liquids [24, 21, 27, 7, 8]. Furthermore, our analysis shows that the spectrum is affected by a wide range of structural properties, as a multitude of features in the ECA curves were non-zero (see Figure 3 and Supplementary Information). In spite of the large number of involved features, we emphasize that this subspace still covers a relatively small fraction of the total structural variance, as indicated in Figures 1b and 1c. Last, the fact that the ECA curves show a contribution from oxygen atoms is a direct indication that the ethanol solvent has a significant effect on the spectrum.
@@
The presented simulations necessitated numerous approximations. For example, the cutoff value for the explicit solvent is an obvious one. We chose the cutoff to be as large as possible while still enabling the simulation of $\sim$104 structure–spectrum pairs. Applying different methods for structural simulation and for property evaluation causes a bias in the results, but the practice is common in spectroscopy and quantum chemistry. In the case of this work, the MD run for the whole simulation cell would not be viable using the electronic structure method of the spectrum calculations. Additionally, we ignore the possible spectral effects arising from vibrational states. Despite this, the reasonability of our data is supported by the fact that the general shape of the spectrum agrees with the experiment. We also note that while switching from the B3LYP to the PBE exhange-correlation potential introduces more transitions (including possible ghost states) and slight changes in the spectra, the discussed structural dependencies of the ROI intensities are consistent across both models.
The challenges and limitations of the ECA method have been discussed in detail in Ref. 8. Most notably, the interpretation of the ECA results is dependent on the properties of the descriptor, intended to encode all information required to predict a spectrum for a structure. While numerous alternatives exist, the LMBTR descriptor excelled in a comparison of six structural descriptor families for X-ray spectra [7], and the performance obtained in the current work for UV–visible aligns well with these earlier results [9, 7, 8]. Further discussion on the required criteria for spectral interpretation by the protocol is available in Ref. 9. In a nutshell, the descriptor must allow good machine learning and ECA performance, and also be interpretable by a human. Development of a new, highly interpretable, atomistic descriptor for machine learning can have a significant positive impact on the method in the future.
@@
Instead of representing two specific atomistic structures, the curves in Figure 3 reflect shifts in the structural distributions between the two selected points in the latent coordinate space. While not necessarily always the case, for the UV–visible spectrum of ethanolic trans-azobenzene, the two latent coordinates link the respective structural classes to a shift of the S2 peak, and reveal potentially important implications for photoprocesses in liquids. Namely, as a consequence of the observed spectral behaviour, the used excitation wavelength introduces a preference for certain initial structures for the following photodynamics, potentially observable by, e.g., pump–probe experiments. Solvent properties have also been found to have an effect on the photochemistry of azobenzene [1], which could partially be a manifestation of this pre-selection.
V Conclusions
@@
The UV–visible absorption spectrum of ethanolic trans-azobenzene shows a broad statistical variation analogous to that found in X-ray spectra. Our subsequent analysis condensed the spectral variance to a few structural degrees of freedom, enabling a partial interpretation of the complex structure–spectrum relationship. We found that photon excitation on the short-wavelength side of the S2 peak favors structures with, among other characteristics, weaker hydrogen bonding with the ethanol solvent, and a contraction of the N=N bond. Thus, the chosen incident photon energy may favor certain structures for the following photodynamics.
Data availability
The data and relevant scripts are available in Zenodo: 10.5281/zenodo.15349624.
Conflicts of interest
There are no conflicts to declare.
@@ Acknowledgements
We acknowledge CSC – IT Center for Science, Finland, for providing computing resources. We thank Dr. N. Runeberg at CSC for support with the Puhti supercomputer, and Prof. J. Wachtveitl for providing the experimental UV–visible absorption spectrum of trans-azobenzene in ethanol solution. J.N. acknowledges funding by Research council of Finland via the academy project grant 367978.
@@
V.1 Validation of the molecular dynamics calculations
To validate the sampling of the molecular dynamics simulations, we evaluated autocorrelation functions of the production run sampled at 10 fs intervals. Figure 5 shows the autocorrelation for ROI intensities, LMBTR vectors, and full spectra. The curves for the ROI values and the full spectra show a rapid initial drop, indicating low correlation even for two temporally subsequent sampled data points. Therefore, the structure–spectrum data set used in this work consists of mostly uncorrelated data points even if the structural autocorrelation decays slower than its spectral counterpart.
Figure 5: Autocorrelation functions of the production run at intervals of 10 fs. a: The curves for z-score standardized data which was used in the analysis. Spectral data is represented by the two regions of interest values (ROIs) and the structural data is represented by the local many-body tensor representation (LMBTR) vectors derived from the xyz-coordinates and atomic number of each nucleus. b Corresponding curves for data that has not been z-score standardized. Additionally, we show the autocorrelation of the full spectra.
V.2 Machine learning details
@@
number of grid points {20, 40}
@@
V.3 A schematic of all LMBTR centers
@@
V.4 ROI value difference with large overall intensity: full list of structural changes
@@
Figure 8 shows the complete list of structural changes implied by a blueshift in the S2 peak, according to the applied analysis procedure. The vector, originally in the z-score-standardized descriptor space, has been inverse transformed to the original descriptor space after which each feature has been divided by the squared distance to the corresponding center. In addition to the results discussed in the main text, the curves at centers C2 and C13 indicate shortening of the para-carbon–meta-carbon bonds. Other less interpretable effects are also implied by the curves; for example, the hydrogen and carbon curves of the virtual centers. We note that the large absolute values near zero exhibited by some of the curves of the virtual centers are an artifact originating from the division by the squared distance to the center.
@@
Figure 8: The full LMBTR vector corresponding to the structural shift behind the increase of the ROI value difference. We denote regions corresponding to weakening of hydrogen bonds with an asterisk (*), shortening of N=N bond with triangles ($\triangledown$), shortening of C–C bonds with diamonds ($\diamond$), and increase in the N–C bond lengths with squares ($\Box$). Numerous other less interpretable factors also play a role in the structure–spectrum relationship.
V.5 Structural changes implied by the increase of total ROI value
@@
Similarly to the analysis of the ROI value difference, Figure 9 shows the complete list of structural changes implied by an increase in the total ROI value, i.e. the sum of the average differential cross section within each ROI. Again, several structural factors affect the spectrum, some being more interpretable than others. The ROI sum is affected by, for example, an increase in the amount of the ethanol solvent near the very tips of the carbon rings. The result can be affected by a molecule being excluded from the explicitly treated cluster due to the necessitated cutoff distance, but is still indicative of a closer ethanol molecule.
@@
Figure 9: The full LMBTR vector corresponding to the structural shift behind the increase in total ROI value. We denote regions corresponding to the increase in the amount of ethanol solvent near the tips of the azobenzene rings with plus signs (+). Many other less interpretable factors also play a role in the structure–spectrum relationship.
V.6 Results using spectra calculated with the PBE exchange-correlation potential
In the main text we studied regions of interest derived from spectra calculated using the B3LYP functional. Here we present the corresponding results using spectra calculated with the PBE exchange-correlation potential. Figure 10 shows the ensemble averaged spectrum, where each line of the individual spectra is convolved with a Gaussian of full-width at half maximum of 0.370 eV and shifted by 0.527 eV. Analogous to the case of B3LYP in the main text, we obtained these values by a fitting the ensemble-averaged spectrum on the experiment by Nägele and coworkers [^4] T. Nägele et. al.: Femtosecond photoisomerization of cis-azobenzene. Chemical Physics Letters 272, 489, 1997. doi:10.1016/S0009-2614(97)00531-9. Especially at the low-wavelength tail of the S2 peak, the simulated spectrum approximates the experiment less well than the corresponding spectrum calculated with B3LYP. The R2 score behavior of the ECA decomposition is similar between the two cases, with the spectral R2 score rising slightly faster with PBE. Figure 11 shows that the first two ECA components show the similar behavior with either of the functionals, except the order of the vectors one and two are flipped. Figure 12 shows that curves denoting structural changes behind a blueshift in the spectrum show highly similar behavior to that found in the main text. Therefore, the qualitative conclusions of the main text are the same for PBE and B3LYP. The different behavior at short distances is likely due to the different LMBTR Gaussian widths originating from the independent model selections performed for B3LYP and PBE data. Figure 13 confirms the consistency in terms of N=N bond lengths and the number of hydrogen bonds.
Figure 10: The corresponding presentation of Figure 1 of the main text, obtained using the PBE functional.
Figure 11: The corresponding presentation of Figure 2 of the main text, obtained using the PBE functional.
Figure 12: The corresponding presentation of Figure 3 of the main text, obtained using the PBE functional.
Figure 13: The corresponding presentation of Figure 4 of the main text, obtained using the PBE functional.